OpenStudy (anonymous):

evaluate the expression e^ln15

jimthompson5910 (jim_thompson5910):

In general e^(ln(x)) = x and ln(e^x) = x

OpenStudy (anonymous):

im confused..

jimthompson5910 (jim_thompson5910):

basically the e^x and ln(x) functions undo each other and cancel each other out

jimthompson5910 (jim_thompson5910):

so e^(ln(2)) = 2 for example

jimthompson5910 (jim_thompson5910):

because the ln(2) gives you some number if you plug this result into e^x, then you'll get 2 back again because the two cancel each other out

OpenStudy (anonymous):

so basically the answers 15 .. jut whatever is right ?

jimthompson5910 (jim_thompson5910):

yep e^(ln(15)) = 15

OpenStudy (anonymous):

whatever a is** thanks

jimthompson5910 (jim_thompson5910):

yw